# spacing=nonfrench

### v3

@@ 1 Introduction
This motivates us to explore a new paradigm for file system design and development, generative file systems, which leverages the capabilities of Large Language Models (LLMs) [chatgpt, liu2024deepseek, liu2024deepseek2, team2024gemini, bai2023qwen] to generate a complete file system and evolve it effectively.
Nevertheless, although recent advancements in LLMs have demonstrated their profound capabilities in automated code generation [dong2025qimeng, cursor, copilot, qi2025intention, zhang2024codeagentenhancingcodegeneration, huang2024agentcodermultiagentbasedcodegeneration, wang2025coderagbenchretrievalaugmentcode, zhang2024autocoderover], generating a complete and useful file system from natural-language prompts is profoundly challenging, if not impossible. The intricate semantics of file systems, from ensuring concurrency correctness to carefully managing complex file structures, are difficult to express unambiguously in prompts. Consequently, to our knowledge, no prior LLM-based approach has successfully generated a complete, functional, and feature-rich file system.
Compared with descriptions with natural language, specifications [stoica2024specificationsmissinglinkmaking, ntzik_et_al:LIPIcs.ECOOP.2018.4] usually provide a precise and machine-understandable language to explicitly encode these invariants (e.g., no orphan inodes after a crash), creating an opportunity to guide LLMs toward generating expected code. However, bridging the gap between specifications and LLM-based code generation is non-trivial. We identify three technical challenges that our work, SysSpec, is designed to overcome:
@@
Challenge II: Complex component composition. The finite context window of LLMs precludes generating an entire file system monolithically. This necessitates a modular approach, which introduces significant composition challenges. First, generating one module at a time, an LLM lacks the global context to ensure interface compatibility with other, yet-to-be-generated or pre-existing components, leading to integration errors. Second, each individual change (with LLMs) introduced during the evolution process can potentially trigger cascading effects on other file system modules, particularly those with existing dependencies. For example, a seemingly local feature addition, like introducing extent [linux-ext4], can trigger non-local changes by altering core data structures like the inode, affecting any module that interacts with it and making manual dependency management intractable.
Challenge III: Unreliable LLM capability. LLM-based code generation is inherently non-deterministic due to the hallucination [10.1145/3703155, xu2025hallucinationinevitableinnatelimitation] — even identical specifications can yield different and potentially incorrect code outputs across generation attempts. A naive “generate-and-pray” approach is unacceptable for system software. Therefore, a robust framework cannot blindly trust the LLM especially for file systems; it must incorporate a rigorous validation mechanism to ensure that generated code strictly adheres to the guiding specification, guaranteeing correctness despite the unreliability.
@@
To address these challenges, our key insight is to replace ambiguous natural language with specifications following the principles adapted from successful practices of formal methods [zou2019atomfs, chen2015using, 199293]. Specifically, we introduce SysSpec, a framework for correct-by-construction file system generation built on three core techniques.
First, we design a formal method-inspired specification to precisely define a file system’s behavior, moving beyond the ambiguity of natural language. It holistically captures the FS design, spanning Functionality, defined through Hoare-logic based pre/post-conditions [zou2019atomfs, chen2015using] and invariants; Modularity, which enforces a clean decomposition into modules with rely-guarantee [zou2019atomfs, Liang12popl] interfaces for correct composition; and Concurrency, which makes locking protocols and ordering explicit to mitigate subtle bugs.
@@
Overall, the new paradigm with SysSpec shifts the developer’s burden from low-level implementation to high-level design. While this demands greater upfront design effort, akin to the safety discipline of Rust [safe-rust, bugden2022rustprogramminglanguagesafety], the return is a file system that is far easier to maintain and evolve.
We utilize SysSpec to implement SpecFS, a complete FUSE-based (generative) file system specified entirely in specification. Our toolchain automatically generates a functional C-language implementation without any manual intervention. We validate its correctness and expressiveness by successfully generating implementations of a previously verified file system, AtomFS [zou2019atomfs]. Besides, we showcase its capability for complex feature integration by seamlessly evolving SpecFS with spec patches to support 10 novel features from Ext4, including delayed allocation and file encryption. We also highlight its benefits on performance: by applying Ext4-style delayed allocation method with a concise spec patch, SysSpec automatically regenerated the relevant modules, resulting in a 99.9% data write reduction for xv6 compilation.
@@ 2.1 The Anatomy of File System Change
Methodology. We choose Ext4 as our target because it is a mature and (still) widely-deployed file system with 20 years of evolution in real-world environments. Ext4 was introduced in Linux 2.6.19 and has been continuously developed, incorporating thousands of patches related to performance, new features, and maintenance. Our analysis is based on all 3,157 Ext4-related commits (i.e., individual patches rather than patch sets) merged into the mainline Linux from version 2.6.19 to 6.15. To understand the evolution, we categorize each commit using a classification scheme for FS patches adapted from prior work [evolutionLu2014]: (1) Bug: fixing an existing bug, (2) Performance: improving efficiency through new designs or optimizations, (3) Reliability: enhancing the file system’s robustness, (4) Feature: implementing new functionalities, and (5) Maintenance: refactoring code or improving documentation without changing semantic behavior.
@@ 2.3 A New Paradigm: Generative File System
0 to NN Copilot [copilot] Natural Language
Clover [sun2024cloverclosedloopverifiablecode] Docstring + Annotation
Qimeng [dong2025qimeng] Programming Language
NN to $N+1$ AutoCodeRover [zhang2024autocoderover] Github Issue
CodeAgent [zhang2024codeagentenhancingcodegeneration] Natural Language
“Intention” [qi2025intention] Half Natural Language
@@
Limitations of prior works. Although abundant prior works achieve automatic code generation, their inherent limitations prevent us from leveraging them to achieve the paradigm for complex file systems, as generalized in Tab. 1. These works can be broadly categorized into two types. First, some works focus on generating complete code logic from scratch (“from 0 to NN”). However, this category of work struggles to cope with the complexity of file systems, whose modules exhibit intricate interactions and dependencies. Consequently, such approaches either produce only simple code (e.g., frontend pages) or are limited to specific single-module tasks [sun2024cloverclosedloopverifiablecode, dong2025qimeng], such as generating tensor program implementations for different hardware [dong2025qimeng], generating unit tests [sotiropoulos2024thalia, chen2022codetcodegenerationgenerated], mathematical library functions [briggs2024megalibm], or performing bit-vector synthesis [ding2024dryadsynth].
Second, other approaches usually focus on evolution, i.e., modifying existing code (“from $N$ to N+1N+1”). They may utilize methods like Retrieval-Augmented Generation (RAG) [wang2025coderagbenchretrievalaugmentcode, qi2025intention] or Agents [cursor, copilot, qi2025intention, zhang2024autocoderover, zhang2024codeagentenhancingcodegeneration, huang2024agentcodermultiagentbasedcodegeneration] to enhance general code generation capabilities. However, these methods largely rely on natural language descriptions (e.g., document strings, GitHub issues [zhang2024autocoderover] or code review comments [qi2025intention]) for the desired program logic. Even though some methods attempt to enhance the LLM’s capability to understand intents expressed in natural language [qi2025intention], ambiguity remains unavoidable: one cannot simply instruct an LLM to “avoid race conditions” and expect a correct outcome. Furthermore, these methods often include all project-related code in the context, requiring the Agent to autonomously decide how to retrieve information from this context. Such a burden is uncontrollable, and an excessively long context can potentially degrade the quality of code generation.
@@ 3 Design Overview
The SysSpec specification. At the core of our framework is a multi-part specification that captures a developer’s design intent, as shown in Fig. 5-a. It consists of: (1) Functionality specifications, which use concepts like Hoare logic (pre/post-conditions) and invariants to describe the behavior of individual modules. (2) Modularity specifications, which decompose the system into distinct components and use a rely-guarantee discipline [10.1145/1480881.1480922] to ensure they can be composed correctly and developed independently. (3) Concurrency specifications, which explicitly define locking protocols and other concurrency-related behaviors that are notoriously difficult for LLMs to infer on their own.
@@ 4.1 Functionality Specification
Hoare logic for file system synthesis. To specify function behavior, we adapt the classic Hoare logic formalism of {P}C{Q} using pre- and post-conditions, but we sidestep the high complexity of full formal verification. Each function is annotated with pre-conditions that define required system states and post-conditions that guarantee specific state transitions and return values. Fig. 6 shows atomfs_ins, an internal function of AtomFS [zou2019atomfs] that implements mknod/mkdir. The pre-condition describes the validness of the parameters. The post-condition describes the state after the operation.
@@ 4.2 Modularity Specification
Context-bounded modular synthesis. One of the core innovations in SysSpec’s modularity specification lies in aligning module design with two intrinsic properties of LLM reasoning. First, strict size constraints ensure each module fits entirely within the model’s context window, enabling holistic analysis during generation. The specific constraints on module sizes evolve in tandem with improvements in LLM capabilities and context window capacities. Take our case study (§5) as an example, we limited module sizes to $\leq$500 LoC, which keeps the token consumption for inference generally within approximately 30K tokens. Second, explicit interface contracts govern inter-module dependencies through Rely-Guarantee conditions, a formal mechanism adapted from concurrent program verification [zou2019atomfs]. This methodology embodies three critical principles: (1) module implementations must respect their declared dependencies (Rely), (2) provide guarantees about their behavior (Guarantee), and (3) compose through logical implication of these contracts.
@@ 5.1 Prototype Implementation
Overview. To demonstrate the benefits of SysSpec, we introduce SpecFS, a concurrent in-memory file system that runs in userspace via FUSE. The architecture is based on a prior formally verified file system, AtomFS [zou2019atomfs]. However, rather than porting its C implementation, we undertake a complete reimplementation guided by our specification. Specifically, we implement AtomFS’s high-level design with SysSpec’s specifications, including pre- and post-conditions (we re-use some conditions from AtomFS’s formal specifications in Coq), invariants, rely/guarantee contracts, system algorithm descriptions, and concurrency specification.
SpecFS is organized into 45 distinct modules distributed across several logical layers, including file operations, inode management, path traversal, and the POSIX interface. SpecFS supports a wide range of standard POSIX calls, such as $\mathsf{open}$, $\mathsf{read}$, and $\mathsf{rename}$. To validate its functional correctness, we use the xfstests [xfstest] suite within our SpecValidator. SpecFS demonstrates a level of correctness equivalent to that of the original AtomFS, failing only 64 out of 754 test cases, all attributable to unimplemented functionality.
@@ 6.1 Evaluations on Accuracy
Methodology. We first study whether SpecFS can accurately generate code to achieve self-evolution, that is, whether the code generation can accurately conform to the logic described by the specification and accurately meet the functionality of the corresponding file system module. Our test cases comprise all the modules to implement the complete logic of AtomFS [zou2019atomfs]. Specifically, to evaluate the accuracy, we first define 45 distinct modules within AtomFS and manually author their ground-truth implementations. A generated module is considered correct if it (i) passes all functional tests for AtomFS and (ii) is deemed logically equivalent to the ground-truth through manual inspection.
We implement two versions of baselines based on a few-shot learning approach. The normal version incorporates a description of the file correspondence logic and the APIs of the dependency modules; The oracle version not only includes the dependency module’s APIs but also integrates the ground-truth code of these modules as part of the prompt. We test generation accuracy using four LLMs of decreasing capability: Gemini-2.5-Pro [gemini], DeepSeek-V3.1 Reasoning [deepseekai2025deepseekv3technicalreport], GPT-5-minimal [gpt5], and Qwen3-32B [yang2025qwen3technicalreport], whose performance is ranked according to the LiveCodeBench leaderboard [livecodebench-leaderboard].
@@ 6.5 Performance Optimizations
Inline data. We first evaluate the compression effect of “Inline Data” on file sizes. As shown in Fig. 13-left, using inline data significantly reduces the file sizes of QEMU [qemu] and Linux source code [linux-src], decreasing the required storage capacity by 35.4% and 21.0%, respectively.
@@ 6.6 Limitations and Discussion
Missing FS features. Although SpecFS represents a successful endeavor in realizing a complex file system using SysSpec, SpecFS is currently implemented as a user-space file system based on FUSE [fuse]. Consequently, it does not operate in kernel mode and lacks a storage stack, such as direct disk access, nor does it consider crash consistency. Our evaluation of SpecFS primarily focuses on validating the generation methodology and ensuring correctness, which precludes an apple-to-apple comparison with native kernel-space file systems regarding raw performance metrics such as throughput.
We plan to apply SpecFS to the self-evolution of industrial file systems (e.g., EROFS or Ext4 [gao2019erofs, linux-ext4]), which is challenging due to the increased engineering complexity. To address this, we consider a fully formally verified “AtomFS-Ext4” unnecessary. Instead, we propose developing a “SpecFS-Ext4” directly based on documentation, using these mitigation strategies. First, instead of specifying the whole state of Ext4, developers can start with a minimal baseline (like Ext2) and incrementally add features to it. Second, we plan to enhance the SpecAssistant to automatically bootstrap draft specifications from documentation (e.g., kernel wikis) or even Ext4 source code. Third, we could adapt methodologies of software engineering to guide the development of specification-constructed (rather than programming-language-constructed) file systems, e.g., applying methods similar to the Law of Demeter [demeterlaw] to reduce the coupling between modules.
@@ 7 Related Work
Domain-specific code generation. Some efforts in program synthesis (e.g., SyGus [sygus]) also focus on automated code generation. They may leverage Domain-Specific Languages (DSLs) to precisely articulate code logic, but face the challenges in scaling to diverse types of scenarios. E.g., MegaLibm [briggs2024megalibm] introduces a novel DSL to construct specifications for mathematical library functions. DryadSynth [ding2024dryadsynth] focuses on bit-vector synthesis building upon the SyGus methodology.
Other research addresses LLM-based code generation for specific domains outside file systems. QiMeng Xpiler [dong2025qimeng] uses LLMs to construct a transcompiler that adapts low-level tensor programs for various hardware platforms. Autoverus [yang2025autoverus] employs LLM-based tools for proving the correctness of Rust code. OSVBench and SpecGen [li2025osvbenchbenchmarkingllmsspecification, ma2025specgenautomatedgenerationformal] address generating formal specifications for a given code snippet.
Formal verification methods. Many studies verify complex systems, such as operating systems [gerwin2009sel4], file systems [zou2019atomfs, zou2024reffs, chen2016fscq], or cloud systems [sun2024anvil]. These efforts also involve formal specifications to ascertain the correctness of handcrafted implementations, but cannot be used diretly for generative file systems. We can draw upon these existing specifications when authoring SysSpec for code generation.
LLM-assisted development. A separate stream of research applies LLMs to mitigate the burden of the file system development. For example, WASABI [stoica2024wasabi] utilizes LLMs to detect intricate retry-related problems in large-scale systems. SysGPT [park2025principles] employs LLMs to provide developers with context-aware performance suggestions. However, they cannot effectively evolve a file system like SpecFS.

### v4

@@ 1 Introduction
This motivates us to explore a new paradigm for file system design and development, generative file systems, which leverages the capabilities of Large Language Models (LLMs) [1, 34, 33, 48, 11] to generate a complete file system and evolve it effectively.
Nevertheless, although recent advancements in LLMs have demonstrated their profound capabilities in automated code generation [20, 3, 2, 24, 53, 25, 49, 54], generating a complete and useful file system from natural-language prompts is profoundly challenging, if not impossible. The intricate semantics of file systems, from ensuring concurrency correctness to carefully managing complex file structures, are difficult to express unambiguously in prompts. Consequently, to our knowledge, no prior LLM-based approach has successfully generated a complete, functional, and feature-rich file system.
Compared with descriptions with natural language, specifications [45, 38] usually provide a precise and machine-understandable language to explicitly encode these invariants (e.g., no orphan inodes after a crash), creating an opportunity to guide LLMs toward generating expected code. However, bridging the gap between specifications and LLM-based code generation is non-trivial. We identify three technical challenges that our work, SysSpec, is designed to overcome:
@@
Challenge II: Complex component composition. The finite context window of LLMs precludes generating an entire file system monolithically. This necessitates a modular approach, which introduces significant composition challenges. First, generating one module at a time, an LLM lacks the global context to ensure interface compatibility with other, yet-to-be-generated or pre-existing components, leading to integration errors. Second, each individual change (with LLMs) introduced during the evolution process can potentially trigger cascading effects on other file system modules, particularly those with existing dependencies. For example, a seemingly local feature addition, like introducing extent [4], can trigger non-local changes by altering core data structures like the inode, affecting any module that interacts with it and making manual dependency management intractable.
Challenge III: Unreliable LLM capability. LLM-based code generation is inherently non-deterministic due to the hallucination [26, 50] — even identical specifications can yield different and potentially incorrect code outputs across generation attempts. A naive “generate-and-pray” approach is unacceptable for system software. Therefore, a robust framework cannot blindly trust the LLM especially for file systems; it must incorporate a rigorous validation mechanism to ensure that generated code strictly adheres to the guiding specification, guaranteeing correctness despite the unreliability.
@@
To address these challenges, our key insight is to replace ambiguous natural language with specifications following the principles adapted from successful practices of formal methods [55, 15, 42]. Specifically, we introduce SysSpec, a framework for correct-by-construction file system generation built on three core techniques.
First, we design a formal method-inspired specification to precisely define a file system’s behavior, moving beyond the ambiguity of natural language. It holistically captures the FS design, spanning Functionality, defined through Hoare-logic based pre/post-conditions [55, 15] and invariants; Modularity, which enforces a clean decomposition into modules with rely-guarantee [55, 31] interfaces for correct composition; and Concurrency, which makes locking protocols and ordering explicit to mitigate subtle bugs.
@@
Overall, the new paradigm with SysSpec shifts the developer’s burden from low-level implementation to high-level design. While this demands greater upfront design effort, akin to the safety discipline of Rust [27, 13], the return is a file system that is far easier to maintain and evolve.
We utilize SysSpec to implement SpecFS, a complete FUSE-based (generative) file system specified entirely in specification. Our toolchain automatically generates a functional C-language implementation without any manual intervention. We validate its correctness and expressiveness by successfully generating implementations of a previously verified file system, AtomFS [55]. Besides, we showcase its capability for complex feature integration by seamlessly evolving SpecFS with spec patches to support 10 novel features from Ext4, including delayed allocation and file encryption. We also highlight its benefits on performance: by applying Ext4-style delayed allocation method with a concise spec patch, SysSpec automatically regenerated the relevant modules, resulting in a 99.9% data write reduction for xv6 compilation.
@@ 2.1 The Anatomy of File System Change
Methodology. We choose Ext4 as our target because it is a mature and (still) widely-deployed file system with 20 years of evolution in real-world environments. Ext4 was introduced in Linux 2.6.19 and has been continuously developed, incorporating thousands of patches related to performance, new features, and maintenance. Our analysis is based on all 3,157 Ext4-related commits (i.e., individual patches rather than patch sets) merged into the mainline Linux from version 2.6.19 to 6.15. To understand the evolution, we categorize each commit using a classification scheme for FS patches adapted from prior work [36]: (1) Bug: fixing an existing bug, (2) Performance: improving efficiency through new designs or optimizations, (3) Reliability: enhancing the file system’s robustness, (4) Feature: implementing new functionalities, and (5) Maintenance: refactoring code or improving documentation without changing semantic behavior.
@@ 2.3 A New Paradigm: Generative File System
0 to NN Copilot [2] Natural Language
Clover [46] Docstring + Annotation
Qimeng [20] Programming Language
NN to $N+1$ AutoCodeRover [54] Github Issue
CodeAgent [53] Natural Language
“Intention” [24] Half Natural Language
@@
Limitations of prior works. Although abundant prior works achieve automatic code generation, their inherent limitations prevent us from leveraging them to achieve the paradigm for complex file systems, as generalized in Tab. 1. These works can be broadly categorized into two types. First, some works focus on generating complete code logic from scratch (“from 0 to NN”). However, this category of work struggles to cope with the complexity of file systems, whose modules exhibit intricate interactions and dependencies. Consequently, such approaches either produce only simple code (e.g., frontend pages) or are limited to specific single-module tasks [46, 20], such as generating tensor program implementations for different hardware [20], generating unit tests [43, 14], mathematical library functions [12], or performing bit-vector synthesis [19].
Second, other approaches usually focus on evolution, i.e., modifying existing code (“from $N$ to N+1N+1”). They may utilize methods like Retrieval-Augmented Generation (RAG) [49, 24] or Agents [3, 2, 24, 54, 53, 25] to enhance general code generation capabilities. However, these methods largely rely on natural language descriptions (e.g., document strings, GitHub issues [54] or code review comments [24]) for the desired program logic. Even though some methods attempt to enhance the LLM’s capability to understand intents expressed in natural language [24], ambiguity remains unavoidable: one cannot simply instruct an LLM to “avoid race conditions” and expect a correct outcome. Furthermore, these methods often include all project-related code in the context, requiring the Agent to autonomously decide how to retrieve information from this context. Such a burden is uncontrollable, and an excessively long context can potentially degrade the quality of code generation.
@@ 3 Design Overview
The SysSpec specification. At the core of our framework is a multi-part specification that captures a developer’s design intent, as shown in Fig. 5-a. It consists of: (1) Functionality specifications, which use concepts like Hoare logic (pre/post-conditions) and invariants to describe the behavior of individual modules. (2) Modularity specifications, which decompose the system into distinct components and use a rely-guarantee discipline [21] to ensure they can be composed correctly and developed independently. (3) Concurrency specifications, which explicitly define locking protocols and other concurrency-related behaviors that are notoriously difficult for LLMs to infer on their own.
@@ 4.1 Functionality Specification
Hoare logic for file system synthesis. To specify function behavior, we adapt the classic Hoare logic formalism of {P}C{Q} using pre- and post-conditions, but we sidestep the high complexity of full formal verification. Each function is annotated with pre-conditions that define required system states and post-conditions that guarantee specific state transitions and return values. Fig. 6 shows atomfs_ins, an internal function of AtomFS [55] that implements mknod/mkdir. The pre-condition describes the validness of the parameters. The post-condition describes the state after the operation.
@@ 4.2 Modularity Specification
Context-bounded modular synthesis. One of the core innovations in SysSpec’s modularity specification lies in aligning module design with two intrinsic properties of LLM reasoning. First, strict size constraints ensure each module fits entirely within the model’s context window, enabling holistic analysis during generation. The specific constraints on module sizes evolve in tandem with improvements in LLM capabilities and context window capacities. Take our case study (§5) as an example, we limited module sizes to $\leq$500 LoC, which keeps the token consumption for inference generally within approximately 30K tokens. Second, explicit interface contracts govern inter-module dependencies through Rely-Guarantee conditions, a formal mechanism adapted from concurrent program verification [55]. This methodology embodies three critical principles: (1) module implementations must respect their declared dependencies (Rely), (2) provide guarantees about their behavior (Guarantee), and (3) compose through logical implication of these contracts.
@@ 5.1 Prototype Implementation
Overview. To demonstrate the benefits of SysSpec, we introduce SpecFS, a concurrent in-memory file system that runs in userspace via FUSE. The architecture is based on a prior formally verified file system, AtomFS [55]. However, rather than porting its C implementation, we undertake a complete reimplementation guided by our specification. Specifically, we implement AtomFS’s high-level design with SysSpec’s specifications, including pre- and post-conditions (we re-use some conditions from AtomFS’s formal specifications in Coq), invariants, rely/guarantee contracts, system algorithm descriptions, and concurrency specification.
SpecFS is organized into 45 distinct modules distributed across several logical layers, including file operations, inode management, path traversal, and the POSIX interface. SpecFS supports a wide range of standard POSIX calls, such as $\mathsf{open}$, $\mathsf{read}$, and $\mathsf{rename}$. To validate its functional correctness, we use the xfstests [17] suite within our SpecValidator. SpecFS demonstrates a level of correctness equivalent to that of the original AtomFS, failing only 64 out of 754 test cases, all attributable to unimplemented functionality.
@@ 6.1 Evaluations on Accuracy
Methodology. We first study whether SpecFS can accurately generate code to achieve self-evolution, that is, whether the code generation can accurately conform to the logic described by the specification and accurately meet the functionality of the corresponding file system module. Our test cases comprise all the modules to implement the complete logic of AtomFS [55]. Specifically, to evaluate the accuracy, we first define 45 distinct modules within AtomFS and manually author their ground-truth implementations. A generated module is considered correct if it (i) passes all functional tests for AtomFS and (ii) is deemed logically equivalent to the ground-truth through manual inspection.
We implement two versions of baselines based on a few-shot learning approach. The normal version incorporates a description of the file correspondence logic and the APIs of the dependency modules; The oracle version not only includes the dependency module’s APIs but also integrates the ground-truth code of these modules as part of the prompt. We test generation accuracy using four LLMs of decreasing capability: Gemini-2.5-Pro [7], DeepSeek-V3.1 Reasoning [18], GPT-5-minimal [8], and Qwen3-32B [51], whose performance is ranked according to the LiveCodeBench leaderboard [9].
@@ 6.5 Performance Optimizations
Inline data. We first evaluate the compression effect of “Inline Data” on file sizes. As shown in Fig. 13-left, using inline data significantly reduces the file sizes of QEMU [10] and Linux source code [6], decreasing the required storage capacity by 35.4% and 21.0%, respectively.
@@ 6.6 Limitations and Discussion
Missing FS features. Although SpecFS represents a successful endeavor in realizing a complex file system using SysSpec, SpecFS is currently implemented as a user-space file system based on FUSE [5]. Consequently, it does not operate in kernel mode and lacks a storage stack, such as direct disk access, nor does it consider crash consistency. Our evaluation of SpecFS primarily focuses on validating the generation methodology and ensuring correctness, which precludes an apple-to-apple comparison with native kernel-space file systems regarding raw performance metrics such as throughput.
We plan to apply SpecFS to the self-evolution of industrial file systems (e.g., EROFS or Ext4 [23, 4]), which is challenging due to the increased engineering complexity. To address this, we consider a fully formally verified “AtomFS-Ext4” unnecessary. Instead, we propose developing a “SpecFS-Ext4” directly based on documentation, using these mitigation strategies. First, instead of specifying the whole state of Ext4, developers can start with a minimal baseline (like Ext2) and incrementally add features to it. Second, we plan to enhance the SpecAssistant to automatically bootstrap draft specifications from documentation (e.g., kernel wikis) or even Ext4 source code. Third, we could adapt methodologies of software engineering to guide the development of specification-constructed (rather than programming-language-constructed) file systems, e.g., applying methods similar to the Law of Demeter [32] to reduce the coupling between modules.
@@ 7 Related Work
Domain-specific code generation. Some efforts in program synthesis (e.g., SyGus [39]) also focus on automated code generation. They may leverage Domain-Specific Languages (DSLs) to precisely articulate code logic, but face the challenges in scaling to diverse types of scenarios. E.g., MegaLibm [12] introduces a novel DSL to construct specifications for mathematical library functions. DryadSynth [19] focuses on bit-vector synthesis building upon the SyGus methodology.
Other research addresses LLM-based code generation for specific domains outside file systems. QiMeng Xpiler [20] uses LLMs to construct a transcompiler that adapts low-level tensor programs for various hardware platforms. Autoverus [52] employs LLM-based tools for proving the correctness of Rust code. OSVBench and SpecGen [30, 37] address generating formal specifications for a given code snippet.
Formal verification methods. Many studies verify complex systems, such as operating systems [28], file systems [55, 56, 16], or cloud systems [47]. These efforts also involve formal specifications to ascertain the correctness of handcrafted implementations, but cannot be used diretly for generative file systems. We can draw upon these existing specifications when authoring SysSpec for code generation.
LLM-assisted development. A separate stream of research applies LLMs to mitigate the burden of the file system development. For example, WASABI [44] utilizes LLMs to detect intricate retry-related problems in large-scale systems. SysGPT [40] employs LLMs to provide developers with context-aware performance suggestions. However, they cannot effectively evolve a file system like SpecFS.